The following are intraocular pressure (mm Hg) values recorded for a sample of 21 elderly subjects: 14.5 13.9 15.0 17.1 12.0 17.5 14.1 12.9 16.9 13.0 17.4 24.2 12.2 14.4 16.0 10.0 18.5 20.8 16.2 15.9 18.6 Can we conclude from these data that the mean of the population from which the sample was drawn is greater than 14? Let α = 0.05. 1. Write the hypotheses, indicate the claim 2. find the critical value t-value 3. calculate the standardized t -value 4. what is the decision
Answer #1
Values ( X ) \Sigma (X_{i} - \bar{X})^{2}
14.5 1.6045
13.9 3.4846
15 0.5878
17.1 1.7777
12 14.188
17.5 3.0043
14.1 2.7779
12.9 8.218
16.9 1.2844
13 7.6546
17.4 2.6677
24.2 71.1205
12.2 12.7213
14.4 1.8679
16 0.0544
10 33.2548
18.5 7.4709
20.8 25.3341
16.2 0.1877
15.9 0.0178
18.6 8.0276
Total 331.1 207.3065

Mean \bar{X} = \Sigma X_{i} / n
\bar{X} = 331.1 / 21 = 15.7667
Standard deviation S_{X} = \sqrt{\Sigma (X_{i} - \bar{X})^{2}/n-1}
S_{X} = \sqrt{ 207.3065 / 21 -1} = 3.2195

To Test :-

H0 :-  \mu = 14

H1 :-  \mu > 14

Test Statistic :-
t = ( \bar{X} - \mu ) / (S /\sqrt{n})
t = ( 15.7667 - 14 ) / ( 3.2195 /\sqrt{ 21 })
t = 2.5147


Test Criteria :-
Reject null hypothesis if t \; > \;t_{\alpha, n-1}
t_{\alpha, n-1} = t_{0.05 , 21-1} = 1.725
t > t_{\alpha, n-1} = 2.5147 > 1.725
Result :- Reject null hypothesis


Decision based on P value
P - value = P ( t > 2.5147 ) = 0.0103
Reject null hypothesis if P value < \alpha = 0.05   level of significance
P - value = 0.0103 < 0.05 ,hence we reject null hypothesis
Conclusion :- Reject null hypothesis

There is sufficient evidence to support the claim that the mean of the population from which the sample was drawn is greater than 14 at α = 0.05.